# Advancing Embodied Agent Security: From Safety Benchmarks to Input Moderation

### v1

@@ Abstract
Embodied agents exhibit immense potential across a multitude of domains, making the assurance of their behavioral safety a fundamental prerequisite for their widespread deployment. However, existing research predominantly concentrates on the security of general large language models, lacking specialized methodologies for establishing safety benchmarks and input moderation tailored to embodied agents. To bridge this gap, this paper introduces a novel input moderation framework, meticulously designed to safeguard embodied agents. This framework encompasses the entire pipeline, including taxonomy definition, dataset curation, moderator architecture, model training, and rigorous evaluation. Notably, we introduce EAsafetyBench, a meticulously crafted safety benchmark engineered to facilitate both the training and stringent assessment of moderators specifically designed for embodied agents. Furthermore, we propose Pinpoint, an innovative prompt-decoupled input moderation scheme that harnesses a masked attention mechanism to effectively isolate and mitigate the influence of functional prompts on moderation tasks. Extensive experiments conducted on diverse benchmark datasets and models validate the feasibility and efficacy of the proposed approach. The results demonstrate that our methodologies achieve an impressive average detection accuracy of 94.58%, surpassing the performance of existing state-of-the-art techniques, alongside an exceptional moderation processing time of merely 0.002 seconds per instance.
@@ 1 Introduction
Embodied agents, which possess the ability to dynamically perceive their environment and make autonomous decisions, represent a promising avenue for integrating artificial intelligence into human life. Corresponding author. This technology exhibits significant potential across diverse fields, such as military operations, flexible manufacturing, and domestic services. However, a fundamental prerequisite for the successful deployment of embodied agents in real-world applications is ensuring the behavioral safety of these intelligent agents.
In general, for embodied agents, large language models (LLMs) often serve as the core component for task planning and decision-making, with their outputs subsequently executed by unmanned systems Vemprala et al. (2024); Singh et al. (2023); Song et al. (2023); Huang et al. (2022). In this process, LLMs generate task plans and action command codes based on the input requirements provided to them. Consequently, the inputs to embodied agents play a pivotal role in determining whether the subsequent actions of the intelligent system will be safe or potentially harmful. This underscores the critical importance to effective input moderation mechanisms to ensure the safety and reliability of these systems.
Existing research on input moderation can be broadly categorized into two approaches: external intervention and internal blocking methods. External intervention typically relies on independent detection models to moderate inputs before they are processed by LLMs. Examples include the Moderation API   [^1] https://platform.openai.com/docs/guides/moderation/overview and Llama-Guard Inan et al. (2023). However, these strategies often suffer from high computational costs and communication overhead, primarily due to their reliance on complex detection models and external API calls. In contrast, internal blocking methods leverage the hidden states generated during model inference to detect and block malicious inputs, as demonstrated by Legilimens Wu et al. (2024) and ToxicDetector Liu et al. (2024). These methods achieve an exceptional balance between effectiveness and efficiency, making them particularly suitable for real-time applications, such as embodied agents.
@@ 3.1 EAsafetyBench-Taxonomy
To address the lack of a unified standard for safety risks in embodied agents, we develop a taxonomy based on public perceptions of risk factors in drone applications Wang et al. (2023); Sabino et al. (2022); Smith et al. (2022); Rubagotti et al. (2022). While our taxonomy does not claim to exhaustively cover all potential risks associated with embodied agents, we believe it is general enough to be easily adapted to other robotic systems. Our work is influenced by state-of-the-art taxonomies, particularly those from Llama Guard Inan et al. (2023) and AEGIS Ghosh et al. (2024). For clarity, we categorize existing taxonomies as ”generative risks” and define our work as ”behavioral risks.”
@@ 3.2 EAsafetyBench-Drone
Instruction Generation We leverage GPT-4 Achiam et al. (2023) as the core tool for efficient and accurate data generation, supplemented by human interaction and supervision.
@@
Augmentation and Filtering Given the generation preferences of a single model, we utilize LLama3.1-70B-Instruct Dubey et al. (2024) to semantically rephrase the candidate dataset generated by GPT-4. The rephrasing process employs two approaches: the first enhances expression diversity by converting between active and passive voice, adjusting tone, and replacing synonyms. The second approach, inspired by the deception attack method in BADROBOT Zhang et al. (2024), subtly replaces concepts to lead embodied agents into performing harmful actions without recognizing their intent. This increases the subtlety of malicious instructions at the semantic level. However, some malicious instructions may lose their harmfulness after rephrasing. To address this, we employ Llama as a judge, utilizing techniques from Zheng et al. (2023) to assess the consistency of rephrased instructions with the originals, specifically regarding malicious intent and harmful consequences. Instructions that deviate significantly from the originals are filtered based on the consistency score.
Dataset Analysis Following the above steps, we construct a safety-aware dataset containing 9,435 instructions, with 4,875 posing safety threats. We present a metric-based analysis of the length distribution and self-BLEU score Zhu et al. (2018) for EAsafetyBench-Drone. The results show that instruction lengths are mostly between 10 and 20 words, with a balanced distribution across categories, and the self-BLEU score is 0.292, indicating a relatively high text diversity.
@@ 3.3 EAsafetyBench-Prompt
We collect 48 prompts from existing research on embodied agents Zawalski et al. (2024); Chen et al. (2023); Huang et al. (2023); TAZIR et al. (2023), covering various agent forms and different formats. These prompts are key components that endow agents with powerful functionalities. Additionally, we gather 52 system prompts related to tasks such as planning or code generation from the open-source repository chatgpt_system_prompt  [^2] https://github.com/LouisShark/chatgpt_system_prompt, enabling a more accurate evaluation of detection methods across diverse prompt contexts.
@@ 4.1 Overview of Pinpoint
When the inference reaches the middle layer (e.g., the 10th layer), the transformer block in this layer performs an additional attention operation based on the instruction mask sequence, generating classification features. The trained classifier then evaluates the instructions, classifying them as either ”malicious” or ”safe.” If malicious, the generation process is immediately halted and returns a predefined refusal response.
@@ 4.3 Intrinsic Feature Extraction
We observe that during the inference, multi-layer attention mechanisms may excessively blend the semantics of instructions and prompts. To address this, we extract features from the middle layers $\mathcal{H}_{middle}$ of the LLMs. In contrast, deep features lead to model overfitting to specific functional prompts, while shallow features are not suitable for classification as they have yet to sufficiently capture the semantic distinctions required for the task, $i.e.$,
$middle=\begin{cases}\text{10,}&\text{if }l\leq 28;\\ \text{17,}&\text{if }l>28.\end{cases}$ (8)
@@ 5.1 Setup
Datasets In addition to EAsafetyBench-Drone, our experiments also utilize data from SafeAgentBench Yin et al. (2024). To the best of our knowledge, these are the only two safety-aware datasets specifically dedicated to embodied agents. SafeAgentBench includes 750 instructions, covering 10 potential hazards: 300 hazardous instructions with safe counterparts, 100 abstract hazardous instructions, and 50 long-horizon instructions with safety risks.
@@
Baselines To evaluate the effectiveness of Pinpoint, we employ three state-of-the-art internal methods as baselines.
@@
Host Model We evaluate the Pinpoint on 10 open-source LLMs. These LLMs encompass various architectures and cover a range of hidden layers from 16 to 40.
@@
Metrics The metrics we adopted include Accuracy, F1 Score, False Negative Rate (FNR), and False Positive Rate (FPR). Accuracy and F1 Score indicate general classification performance. FNR measures the proportion of malicious instructions misclassified as safe, which could pose safety risks. FPR measures the proportion of safe instructions misclassified as malicious, which could impact user experience.
Experimental Settings All experiments are conducted on Ubuntu 22.04 using four NVIDIA RTX A6000 GPUs. We build the entire experimental environment on the PyTorch platform and configure all baselines and host LLMs strictly following their respective specifications. We train a fully connected MLP classifier with 3 layers and 4 million parameters using the Adam optimizer. The training parameters are set as follows: a batch size of 16, 50 epochs, a learning rate of 1e-3, and a weight decay ($\ell_{2}$ penalty) of 2e-4.
@@
5.2 Performance Evaluation
Effectiveness In this section, we evaluate the effectiveness of Pinpoint. We propose two strategies to select the layer m for feature extraction: for LLMs with 28 or fewer layers, m=10, and for LLMs with more than 28 layers, m=17. To comprehensively validate Pinpoint’s performance, we compare its results on multiple host LLMs against baselines. Experimental results are presented in Table 1.
@@ 5.2 Performance Evaluation
Notably, Pinpoint exhibits relatively low metrics on the ChatGLM which is based on a prefix decoder architecture. This is primarily because the prefix decoder employs a bidirectional attention mechanism when processing inputs. Unlike causal mask, it incorporates more semantic information from the functional prompt, particularly when the instructions are embedded in the middle of the prompt. This issue can be addressed by concatenating multi-layer features.
Efficiency In this section, we measure the average time overhead for different methods in completing detection tasks on the test set which is embedded within in-the-wild prompts. To simulate real-world applications, we set the batch size to 1 for all tests. The measurement results are shown in Table 4. Compared to the baselines, Pinpoint outperforms all other methods in terms of efficiency, requiring only 0.002 seconds per moderation. Furthermore, when deployed on different host LLMs, Pinpoint maintains the same level of efficiency, meeting the real-time requirements of embodied agents.
@@ 5.3 Ablation Study
Layer Selection Strategy Figure 3 illustrates the performance of LLMs with varying numbers of layers in feature extraction for detection at each layer. The trends in each subplot are consistent, with performance gradually declining or stabilizing starting from the middle layers. This aligns with our previous analysis: deeper layers integrate more semantics of functional prompts through attention mechanisms, which may introduce additional noise and impact detection performance. This effect is particularly pronounced in ChatGLM, which is based on a prefix decoder architecture.
Instruction Masked Attention In Figure 3, the orange and gray curves represent classification performance on visible and in-the-wild prompts, respectively, using a method without masked attention (directly using the final layer’s last token hidden state as classifier input). Across all host LLMs, incorporating masked attention significantly improves performance, particularly in generalizing to in-the-wild prompts. By leveraging this, Pinpoint effectively handles the complex and diverse prompts encountered in real-world applications.

### v3

@@ Abstract
Embodied agents exhibit immense potential across a multitude of domains, making the assurance of their behavioral safety a fundamental prerequisite for their widespread deployment. However, existing research predominantly concentrates on the security of general large language models, lacking specialized methodologies for establishing safety benchmarks and input moderation tailored to embodied agents. To bridge this gap, this paper introduces a novel input moderation framework, meticulously designed to safeguard embodied agents. This framework encompasses the entire pipeline, including taxonomy definition, dataset curation, moderator architecture, model training, and rigorous evaluation. Notably, we introduce EAsafetyBench, a meticulously crafted safety benchmark engineered to facilitate both the training and stringent assessment of moderators specifically designed for embodied agents. Furthermore, we propose Pinpoint, an innovative prompt-decoupled input moderation scheme that harnesses a masked attention mechanism to effectively isolate and mitigate the influence of functional prompts on moderation tasks. Extensive experiments conducted on diverse benchmark datasets and models validate the feasibility and efficacy of the proposed approach. The results demonstrate that our methodologies achieve an impressive average detection accuracy of 94.58%, surpassing the performance of existing state-of-the-art techniques, alongside an exceptional moderation processing time of merely 0.002 seconds per instance. The source code and datasets can be found at https://github.com/ZihanYan-CQU
/EAsafetyBench
@@ 1 Introduction
Embodied agents, which possess the ability to dynamically perceive their environment and make autonomous decisions, represent a promising avenue for integrating artificial intelligence into human life.
This technology exhibits significant potential across diverse fields, such as military operations and domestic services. However, a fundamental prerequisite for the successful deployment of embodied agents in real-world applications is ensuring the behavioral safety of these intelligent agents.
In general, for embodied agents, large language models (LLMs) often serve as the core component for task planning and decision-making, with their outputs subsequently executed by unmanned systems Vemprala et al. (2024); Singh et al. (2023); Song et al. (2023); Huang et al. (2022). In this process, LLMs generate task plans and action command codes based on the input requirements provided to them. Consequently, the inputs to embodied agents play a pivotal role in determining whether the subsequent actions of the intelligent system will be safe or potentially harmful. This underscores the critical importance of effective input moderation mechanisms for ensuring the safety and reliability of these systems.
Existing research on input moderation can be broadly categorized into two approaches: external intervention and internal blocking. External approaches typically employ standalone detection models like Moderation API [^1] https://platform.openai.com/docs/guides/moderation/overview and Llama Guard Inan et al. (2023) to filter input before LLM processing, but incur high computational costs due to complex detection models and external API calls. Instead, internal blocking methods leverage the hidden states generated during model inference to detect and block malicious input, as demonstrated by Legilimens Wu et al. (2024) and ToxicDetector Liu et al. (2024). These methods strike an exceptional balance between effectiveness and efficiency, making them ideal for real-time applications, such as embodied agents.
@@ 3.1 EAsafetyBench-Taxonomy
To address the lack of a unified standard for safety risks in embodied agents, we develop a taxonomy based on public perceptions of risk factors in drone applications Wang et al. (2023); Sabino et al. (2022); Smith et al. (2022); Rubagotti et al. (2022). While our taxonomy does not claim to exhaustively cover all potential risks associated with embodied agents, we believe it is general enough to be easily adapted to other robotic systems. Our work is influenced by state-of-the-art taxonomies, particularly those from Llama Guard Inan et al. (2023) and AEGIS Ghosh et al. (2024). For clarity, we categorize existing taxonomies as ”generative risks” and define our work as ”behavioral risks”.
@@ 3.2 EAsafetyBench-Drone
Instruction Generation
We leverage GPT-4 Achiam et al. (2023) as the core tool for efficient and accurate data generation, supplemented by human interaction and supervision.
@@
Augmentation and Filtering
Given the generation preferences of a single model, we utilize LLama3.1-70B-Instruct Dubey et al. (2024) to semantically rephrase the candidate dataset generated by GPT-4. The rephrasing process employs two approaches: the first enhances expression diversity by converting between active and passive voice, adjusting tone, and replacing synonyms. The second approach, inspired by the deception attack method in BADROBOT Zhang et al. (2024), subtly replaces concepts to lead embodied agents into performing harmful actions without recognizing their intent. This increases the subtlety of malicious instructions at the semantic level. However, some malicious instructions may lose their harmfulness after rephrasing. To address this, we employ Llama as a judge, utilizing techniques from Zheng et al. (2023) to assess the consistency of rephrased instructions with the originals, specifically regarding malicious intent and harmful consequences. Instructions that deviate significantly from the originals are filtered based on the consistency score.
Dataset Analysis
Following the above steps, we construct a safety-aware dataset containing 9,435 instructions, with 4,875 posing safety threats. We present a metric-based analysis of the length distribution and self-BLEU score Zhu et al. (2018) for EAsafetyBench-Drone. The results show that instruction lengths are mostly between 10 and 20 words, with a balanced distribution across categories, and the self-BLEU score is 0.292, indicating a relatively high text diversity.
@@ 3.3 EAsafetyBench-Prompt
We collect 48 prompts from existing research on embodied agents Zawalski et al. (2024); Chen et al. (2023); Huang et al. (2023); TAZIR et al. (2023), covering various agent forms and different formats. These prompts are key components that endow embodied agents with powerful functionalities. Additionally, we gather 52 system prompts related to tasks such as planning or code generation from the open-source repository PromptCraft [^2] https://github.com/LouisShark/chatgpt_system_prompt, enabling a more accurate evaluation of detection methods across diverse prompt contexts.
@@ 4.1 Overview of Pinpoint
When the inference reaches the middle layer (e.g., the 10th layer), the transformer block in this layer performs an additional attention operation based on the instruction mask sequence, generating classification features. The trained classifier then evaluates the instructions, classifying them as either ”malicious” or ”safe”. If malicious, the generation process is immediately halted and returns a predefined refusal response.
@@ 4.3 Intrinsic Feature Extraction
We observe that during the inference, multi-layer attention mechanisms may excessively blend the semantics of instructions and prompts. To address this, we extract features from the middle layers $\mathcal{H}_{middle}$ of the LLMs. In contrast, deep features lead to model overfitting to specific functional prompts, while shallow features are not suitable for classification as they have yet to sufficiently capture the semantic distinctions required for the task.
@@ 5.1 Setup
Datasets
In addition to EAsafetyBench-Drone, our experiments also utilize data from SafeAgentBench Yin et al. (2024). To the best of our knowledge, these are the only two safety-aware datasets specifically dedicated to embodied agents. SafeAgentBench focuses on potential hazards to humans or property in household scenarios. It comprises 750 instructions, including 450 tasks with safety risks.
@@
Baselines
To evaluate the effectiveness of Pinpoint, we employ three state-of-the-art internal methods as baselines.
@@
Host Model
We evaluate the Pinpoint on 10 open-source LLMs. These LLMs encompass various architectures and cover a range of hidden layers from 16 to 40.
@@
Metrics
The metrics we adopted include Accuracy, F1 Score, False Negative Rate (FNR), and False Positive Rate (FPR). Accuracy and F1 Score indicate general classification performance. FNR measures the proportion of malicious instructions misclassified as safe, which could pose safety risks. FPR measures the proportion of safe instructions misclassified as malicious, which could impact user experience.
Experimental Settings
All experiments are conducted on Ubuntu 22.04 using four NVIDIA RTX A6000 GPUs. We build the entire experimental environment on the PyTorch platform and configure all baselines and host LLMs strictly following their respective specifications. We train a fully connected MLP classifier with 3 layers and 4 million parameters using the Adam optimizer. The training parameters are set as follows: a batch size of 16, 50 epochs, a learning rate of 1e-3, and a weight decay ($\ell_{2}$ penalty) of 2e-4.
5.2 Performance Evaluation
Effectiveness
In this subsection, we evaluate the effectiveness of Pinpoint. We select the layer m exhibiting the highest overall performance for the extraction of features to perform the experimental presentation. Specifically, for LLMs with 28 or fewer layers, m=10, and for LLMs with more than 28 layers, m=17. To comprehensively validate Pinpoint’s perfor-
@@ 5.2 Performance Evaluation
mance, we compare its results on multiple host LLMs against baselines. Experimental results are presented in Table 1.
@@
Notably, Pinpoint exhibits relatively low metrics on the ChatGLM which is based on a prefix decoder architecture. This primarily stems from the bidirectional attention mechanism in prefix decoders, which inherently incorporates more semantic information from functional prompts. We preliminarily investigate multilayer feature fusion as a viable solution. Specifically, we extract and fuse features from multiple hidden layers (e.g., concatenating features from layers 9, 10, and 11 for LLMs with 28 or fewer layers, concatenating features from layers 16, 17, and 18 for LLMs with more than 28 layers). This approach enables Pinpoint to achieve an F1 score of 0.9498 on prefix-decoder architecture models, which is comparable to that of the causal architecture model.
Efficiency
In this subsection, we measure the average time overhead for different methods in completing detection tasks on the test set which is embedded within in-the-wild prompts. To simulate real-world applications, we set the batch size to 1 for all tests. The measurement results are shown in Table 4. Compared to the baselines, Pinpoint outperforms all other methods in terms of efficiency, requiring only 0.002 seconds per moderation. Furthermore, when deployed on different host LLMs, Pinpoint maintains the same level of efficiency, meeting the real-time requirements of the embodied agents.
@@ 5.3 Ablation Study
Layer Selection Strategy
Figure 3 illustrates the performance of LLMs with varying numbers of layers in feature extraction for detection at each layer. The trends in each subplot are consistent, with performance gradually declining or stabilizing starting from the middle layers. This aligns with our previous analysis: deeper layers integrate more semantics of functional prompts through attention mechanisms, which may introduce additional noise and impact detection performance. This effect is particularly pronounced in ChatGLM, which is based on a prefix decoder architecture.
Instruction Masked Attention
In Figure 3, the orange and gray curves represent classification performance on visible and in-the-wild prompts, respectively, using a method without masked attention (directly using the final layer’s last token hidden state as classifier input). Across all host LLMs, incorporating masked attention significantly improves performance, particularly in generalizing to in-the-wild prompts. By leveraging this, Pinpoint effectively handles the complex and diverse prompts encountered in real-world applications.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant 2022YFB3103500, in part by the National Natural Science Foundation of China under Grant 62101079, in part by the Venture and Innovation Support Program for Chongqing Overseas Returnees under Grant cx2021012.